OpenStudy (anonymous):

What is the equation in slope-intercept form of the line through the points (2, -3) and (-3, 7)

OpenStudy (anonymous):

First, find the slope using \[(y _{2} - y _{1})/(x _{2}-x _{1})\] (7-(-3))/(-3-2) = 7+3/-5=10/-5 = -2 Then, plug the slope and either point into slope intercept form to find b, the y-intercept. y=mx+b -3=-2(2) + b -3=-4+b 1=b Now, you just plug m, the slope, and b, the y-intercept into the equation. y=-2x+1